A duopoly market in which demand is given by 180-2q. The total cost for firm 1 is TC = 60q1, the total cost for firm 2 is TC1 = 66q2. The good is homogeneous.

A) Solve for when Firm 1 chooses the quantity before firm 2. What is the subgame Nash equilibrium?

b) Solve when the firms compete simultaneously. What is the Nash equilibrium?
Answer #1

= P = 180-29, T6,=604, TC 2 = 6692 A sequential choice game . fl chooses first, f2 then then Ta= (P-MC2) 92 = (180- 24-292-66